OpenStudy (anonymous):

a piece of string is 3 yards long how many 1 1/4 yard long pieces can julie cut from the string a 1 b 2 c 3 d 4

OpenStudy (anonymous):

Two

hero (hero):

\[3 \div \left(1 + \frac{1}{4}\right)\]

hero (hero):

Well, that's the hard way to do it. The easy way is to keep adding \(1 + \frac{1}{4}\) to itself

hero (hero):

If you add it to itself you get \(2 + \frac{1}{2}\)

hero (hero):

So that lets you know you can only do it two times without going over 3.

hero (hero):